OpenStudy (ja1):

What are the solutions of 6x^2 = 12x?

OpenStudy (ja1):

x = 0, x = −2 x = 0, x = 2 x = 6, x = 2 x = 6, x = −2

OpenStudy (ja1):

I am brain dead on this one xD I forgot how to do these problems

OpenStudy (anonymous):

Is this supposed to be: \[x^2=12x\]

OpenStudy (ja1):

Yupp

OpenStudy (ja1):

Wait no

OpenStudy (anonymous):

\[6x^2=12x\]

OpenStudy (ja1):

\[6x^{2} = 12x\]

OpenStudy (ja1):

Yah lol

OpenStudy (anonymous):

Ok, start by dividing both sides by 6, so you get this: \[x^2=2x\] Now subtract the 2x from both sides: \[x^2 -2x=0\] From here we can factor out an x and we get: \[x(x-2)=0\] After setting both factors equal to 0 we find x=0 or x=2

OpenStudy (ja1):

Yes but how do you get 2 solutions that's were I am stuck :/

OpenStudy (anonymous):

Once you have the two factors, you set each factor equal to 0. So, we set x=0 and x-2=0. Solving for x on the second one gives us x=2. Try plugging both solutions into the original to check them.

OpenStudy (ja1):

OHHH ok I understand now

OpenStudy (ja1):

I went kinda blank on this one

OpenStudy (ja1):

Thanks :D

OpenStudy (anonymous):

That happens sometimes. Glad I could help.

OpenStudy (ja1):

:D thanks